OpenStudy (anonymous):

write the slope-intercept form of the equations of the line described. 1.through (1,3), parallel to y=5x-5 2.through (1,2), parallel to y=4x+3 3.through(-4,4) perpendicular to y=4/3x+2 4.through (5,5) perpendicular to y=-x+3

OpenStudy (anonymous):

1. y=5x+3 2.y=4x+2 3.y=-3/4x+4 4.y=x+5